Calculate van der Waals' constants for
helium using the data Tc = 5.3 K,
pc = 2.25 atm and R = 8.31 J morl K-1.
1
Expert's answer
2019-05-06T10:36:45-0400

The van der Waals' constants a and b are related to the critical temperature "{{T}_{c}}" and pressure "{{p}_{c}}" by

"a=\frac{27T_{c}^{2}{{R}^{2}}}{64{{p}_{c}}},\,\,\,\,b=\frac{{{T}_{c}}R}{8{{p}_{c}}}"

We have for helium "{{T}_{c}}=5.3\,K" and "{{p}_{c}}\text{ }=\text{ }2.25\text{ }atm." Convert pressure in atmospheres to pressure in pascals:

"{{p}_{c}}=\text{1}\text{.01325}\cdot \text{1}{{\text{0}}^{5}}\,\left( Pa/1\,atm \right)\cdot 2.25\text{ }atm\approx \text{2}\text{.28}\cdot \text{1}{{\text{0}}^{5}}\text{ }Pa"

Substituting the known values into these relations, we get


"a=\frac{27\cdot {{\left( 5.3\,K \right)}^{2}}{{\left( 8.31\text{ }J\cdot mo{{l}^{-1}}\,\cdot {{K}^{-1}} \right)}^{2}}}{64\cdot \text{2}\text{.28}\cdot \text{1}{{\text{0}}^{5}}\text{ }Pa}"

"\approx 0.0036\,\frac{{{J}^{2}}}{Pa\cdot mo{{l}^{2}}}=0.0036\,\frac{Pa\cdot {{m}^{6}}}{mo{{l}^{2}}}"

"b=\frac{5.3\,K\cdot 8.31\text{ }J\cdot mo{{l}^{-1}}\,\cdot {{K}^{-1}}}{8\cdot \text{2}\text{.28}\cdot \text{1}{{\text{0}}^{5}}\text{ }Pa}\approx \text{2}\text{.4}\cdot \text{1}{{\text{0}}^{-5}}\frac{J}{Pa\cdot mol}=\text{2}\text{.4}\cdot \text{1}{{\text{0}}^{-5}}\frac{{{m}^{3}}}{mol}"

To convert a from "Pa\text{ }\cdot {{m}^{6}}/mo{{l}^{2}}" to "Pa\cdot {{L}^{2}}/mo{{l}^{2}}" multiply by 106

"a=3.6\,\cdot {{10}^{3}}\frac{Pa\cdot {{L}^{2}}}{mo{{l}^{2}}}"

To convert b from "{{m}^{3}}\/mo{{l}}" to "L\/mo{{l}}" multiply by 103

"b=\text{2}\text{.4}\cdot \text{1}{{\text{0}}^{-2}}\frac{L}{mol}"

To convert a from "Pa\cdot \text{ }{{L}^{2}}/mo{{l}^{2}}" , to "bar\cdot \text{ }{{L}^{2}}/mo{{l}^{2}}" multiply by 10-5

"a=3.6\,\cdot {{10}^{-2}}\frac{bar\cdot {{L}^{2}}}{mo{{l}^{2}}}"

Thus

"a=0.0036\,\frac{Pa\cdot {{m}^{6}}}{mo{{l}^{2}}}=3.6\,\cdot {{10}^{3}}\frac{Pa\cdot {{L}^{2}}}{mo{{l}^{2}}}=3.6\,\cdot {{10}^{-2}}\frac{bar\cdot {{L}^{2}}}{mo{{l}^{2}}}"

"b=\text{2}\text{.4}\cdot \text{1}{{\text{0}}^{-5}}\frac{{{m}^{3}}}{mol}=\text{2}\text{.4}\cdot \text{1}{{\text{0}}^{-2}}\frac{L}{mol}"